102.     Brother Bear Corporation’s unadjusted trial balance includes the following balances (assume normal balances):

Accounts Receivable                                   $646,000

Allowance for Doubtful Accounts                    15,200

Bad debts are estimated to be 6% of outstanding receivables. What amount of bad debts expense will the company record?

103.     Manning Retailers accepted $75,000 of Citibank Visa credit card charges for merchandise sold on July 1. Citibank charges 4% for its credit card use. The entry to record this transaction by Manning Retailers will include a credit to Sales of $75,000 and a debit(s) to

a.   Cash $72,000 and Service Charge Expense $3,000.

b.   Accounts Receivable $72,000 and Service Charge Expense $3,000.

c.   Cash $72,000 and Interest Expense $3,000.

d.   Accounts Receivable $75,000.

104.     ABC Company accepted a national credit card for a $3,000 purchase. The cost of the goods sold is $2,400. The credit card company charges a 3% fee. What is the impact of this transaction on net operating income?

a.   Increase by $582

b.   Increase by $600

c.   Increase by $510

d.   Increase by $2,910

116.     A 60-day note receivable dated June 13 has a maturity date of

a.   August 13.

b.   August 12.

c.   August 11.

d.   August 10.

117.     The maturity value of a $90,000, 10%, 60-day note receivable dated July 3 is

a.   $90,000.

b.   $99,000.

c.   $105,000.

d.   $91,500.
Answer #1

102. Bad debts are 6%= 6%*646000=$38760

Allowance for bad debts that appear in Financial statements =$15200

So the amount of bad debt expense that shall be recorded =$38760-$15200=$ 23560

103.

A. As the credit card use shall be equal to the cash received, as the manning retailers is being charged for such services by bank. And the bank account of retailer is linked with the bank account hence, cash shall be realised immediately.

104.

Net sale =3000-3%=$2910

Cost of Goods Sold =$2400

Net incom=$510

So the correct answer is C

116. Correct answer is B

August 12

117.

Interest :90000*60/360*10%=$1500

So maturity value is =$90000+$1500

=$91500

Correct answer is D

Do give your feedback! Happy learning :) :)